OpenStudy (calculusxy):

I just did a problem on point-slope form and when I was done with a question, I looked at the answer key for confirmation. At some points, I saw that the answer key had a much simplified version (y + 2 = 4x), whereas there are other times when the answer key said for me to leave it as like y + 5 = -(x - 4). Why is this the case?

OpenStudy (calculusxy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'd have to see the full problem

OpenStudy (calculusxy):

Okay so the problems are: through (0, -2) and (-1, 2) for the answer of y + 2 = 4x through (4, -5) and (0, -1) for the answer of y + 5 = -(x - 4)

jimthompson5910 (jim_thompson5910):

The slope of the line through (0, -2) and (-1, 2) should be negative

OpenStudy (calculusxy):

oh yes sorry y + 2 = -4x

jimthompson5910 (jim_thompson5910):

good, which is the same as `y+2 = -4(x-0)`

OpenStudy (calculusxy):

yes

jimthompson5910 (jim_thompson5910):

they decided to simplify the `x-0` and replace it with just `x`

jimthompson5910 (jim_thompson5910):

the slope of the line through (4, -5) and (0, -1) is not -1

OpenStudy (calculusxy):

why not?

jimthompson5910 (jim_thompson5910):

oh my bad, I did -1-5 instead of -1-(-5). Yes the slope is definitely -1

OpenStudy (calculusxy):

ok

OpenStudy (calculusxy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what other questions did you have?

OpenStudy (phi):

point-slope is y - y0= m(x- x0) where m is the slope and (x0, y0) is the point however , if either x0 or y0 is zero you would simplify (*nobody* bothers to show add or subtract zero) that is why y + 2 = -4(x-0) is written as y+2 = -4x